# On the Learn-to-Optimize Capabilities of Transformers in In-Context Sparse Recovery

### v1

@@ C.1 Transformer Structure
where the submatrix $\mathbf{W}_{1,\text{sub}}$ is defined as
$\mathbf{W}_{1,\text{sub}}=\begin{bmatrix}\mathbf{0}_{(d+1)\times(d+1)}&\mathbf{0}_{(d+1)\times d}&\mathbf{0}_{d+1}\\ \mathbf{0}_{d\times(d+1)}&\mathbf{I}_{d\times d}&\mathbf{0}_{d}\\ \mathbf{0}_{1\times(d+1)}&\mathbf{0}_{1\times d}&0\end{bmatrix}.$
Therefore, we obtain that the output of the MLP layer is
@@
and $|\bm{\beta}_{2n+1}^{(k-1)}|\leq C_{\bm{\beta}}^{k}$. Therefore, we have
$|\bm{\beta}_{i}^{(k)}|$ $\leq|\bm{\beta}_{i}^{(k-1)}|+|\gamma^{(k-1)}|\|\mathbf{M}^{V}\|\left\|\frac{1}{i}[\mathbf{X}]_{1:\lfloor\frac{i+1}{2}\rfloor,:}^{\top}[\mathbf{X}]_{1:\lfloor\frac{i+1}{2}\rfloor,:}\right\||\bm{\beta}_{i}^{(k-1)}|+b_{\bm{\beta}}|\gamma^{(k-1)}|\|\mathbf{M}^{V}\|\left\|\frac{1}{i}[\mathbf{X}]_{1:\lfloor\frac{i+1}{2}\rfloor,:}\right\|$
@@
Equation C.9 arises from the fact that $\lfloor\frac{i+1}{2}\rfloor/i\leq 1$. Therefore, denoting $\gamma_{\max}=\max_{k}\|\gamma^{(k)}\|$ and letting $C_{\bm{\beta}}=1+b_{\mathbf{x}}^{2}d\gamma_{\max}\|\mathbf{M}^{V}\|+b_{\bm{\beta}}b_{\mathbf{x}}\gamma_{\max}\|\mathbf{M}^{V}\|$, we have
@@
$\sum_{u\in\{-1,+1\}}\sum_{j=1}^{i}\sigma\left(\Big{\langle}\mathbf{Q}^{(k)}_{u}\mathbf{h}_{i}^{k},\mathbf{K}^{(k)}_{u}\mathbf{h}_{j}^{k}\Big{\rangle}\right)\cdot\mathbf{V}_{u}^{(k)}\mathbf{h}^{(k)}_{j}$ $=\sum_{u\in\{-1,+1\}}\sum_{j=2}^{i-1}\sigma\left(\Big{\langle}\mathbf{Q}^{(k)}_{u}\mathbf{h}_{i}^{k},\mathbf{K}^{(k)}_{u}\mathbf{h}_{j}^{k}\Big{\rangle}\right)\cdot\mathbf{V}_{u}^{(k)}\mathbf{h}^{(k)}_{j}$
@@
$\sum_{u\in\{-2,+2\}}\sum_{j=1}^{i}\sigma\left(\Big{\langle}\mathbf{Q}^{(k)}_{u}\mathbf{h}_{i}^{k},\mathbf{K}^{(k)}_{u}\mathbf{h}_{j}^{k}\Big{\rangle}\right)\cdot\mathbf{V}_{u}^{(k)}\mathbf{h}^{(k)}_{j}$ $=\sum_{u\in\{-2,+2\}}\sum_{j=2}^{i-1}\sigma\left(\Big{\langle}\mathbf{Q}^{(k)}_{u}\mathbf{h}_{i}^{k},\mathbf{K}^{(k)}_{u}\mathbf{h}_{j}^{k}\Big{\rangle}\right)\cdot\mathbf{V}_{u}^{(k)}\mathbf{h}^{(k)}_{j}$
@@
$\mathbf{h}_{i}^{(k+1)}$ $=\mathrm{MLP}\left(\mathbf{h}_{i}^{(k)}+\frac{1}{i}\sum_{u\in\{\pm 1,\pm 2\}}\sum_{j=1}^{i}\sigma\left(\Big{\langle}\mathbf{Q}^{(k)}_{u}\mathbf{h}_{i}^{k},\mathbf{K}^{(k)}_{u}\mathbf{h}_{j}^{k}\Big{\rangle}\right)\cdot\mathbf{V}_{u}^{(k)}\mathbf{h}^{(k)}_{j}\right)=\begin{bmatrix}[\mathbf{X}^{\top}]_{:,\lfloor\frac{i+1}{2}\rfloor}\\ 0\\ \bm{\beta}^{k+1}_{i}\\ 1\end{bmatrix},$
@@
$\mathbf{h}_{i}^{(k+1)}$ $=\mathrm{MLP}\left(\mathbf{h}_{i}^{(k+1)}+\frac{1}{i}\sum_{u\in\{\pm 1,\pm 2\}}\sum_{j=1}^{i}\sigma\left(\Big{\langle}\mathbf{Q}^{(k)}_{u}\mathbf{h}_{i}^{k},\mathbf{K}^{(k)}_{u}\mathbf{h}_{j}^{k}\Big{\rangle}\right)\cdot\mathbf{V}_{u}^{(k)}\mathbf{h}^{(k)}_{j}\right)=\begin{bmatrix}[\mathbf{X}^{\top}]_{:,\lfloor\frac{i+1}{2}\rfloor}\\ 0\\ \bm{\beta}^{k+1}_{i}\\ 0\end{bmatrix}.$ (C.13)
@@
Recall that $c_{2}=\log\left(\frac{1}{\gamma^{(k)}\sigma_{\max}(2S-1)+\left|1-\gamma^{(k)}\sigma_{\min}\right|}\right)$, we obtain
@@
Next, we introduce Proposition 2.
@@
$\gamma\leq\min\left\{\frac{1}{\sigma_{\min}},\frac{1}{\max\{([\mathbf{D}_{n}]_{:,i})^{\top}\mathbf{X}_{1:n,i}\}}\right\},$ (D.11)
then Inequality D.6 will hold and $1-\gamma\sigma_{\min}\geq 0$. Besides, we have
@@
we have $\gamma$ satisfies Equation D.11. Note that we assume $n\geq N_{0}$. Therefore,
@@
$\|\bm{\beta}^{(k)}_{2n+1}-\bm{\beta}^{*}\|$ $\leq\|\bm{\beta}^{*}\|_{1}\Big{(}\gamma(2S-1)\sigma_{\max}+\left|1-\gamma\sigma_{\min}\right|\Big{)}^{k-1}$
@@ D.2 Proof of Remark 3
Therefore, the output of the modified MLP layer is
@@
the corresponding output of the self-attention layer satisfies
@@
Therefore, the proof is complete. ∎
@@
$\frac{1}{i}\sum_{j=1}^{i}\left\langle\mathbf{Q}_{1}^{\mathrm{last}}\mathbf{h}_{i}^{(K+1)},\mathbf{K}_{1}^{\mathrm{last}}\mathbf{h}_{j}^{(K+1)}\right\rangle\cdot\mathbf{V}_{1}^{\mathrm{last}}\mathbf{h}_{j}^{(K+1)}$ $=\frac{1}{i}\sum_{j=1}^{i}\sigma\left(\mathbf{x}_{n}^{\top}\bm{\beta}^{(K+1)}_{j}-B^{\prime}\right)\cdot\begin{bmatrix}\mathbf{0}_{d}\\ \mathbbm{1}_{\{i\%2=1\}}\\ \mathbf{0}_{d+1}\end{bmatrix}$
@@
$\mathrm{MLP}\left(\frac{1}{i}\sum_{j=1}^{i}\left\langle\mathbf{Q}_{1}^{\mathrm{last}}\mathbf{h}_{i}^{(K+1)},\mathbf{K}_{1}^{\mathrm{last}}\mathbf{h}_{j}^{(K+1)}\right\rangle\cdot\mathbf{V}_{1}^{\mathrm{last}}\mathbf{h}_{j}^{(K+1)}\right)=\begin{bmatrix}\mathbf{0}_{d}\\ \frac{2}{2np-p}\sum_{j:o_{j}=1}\mathbf{x}_{n}^{\top}\bm{\beta}^{(K+1)}_{j}-\frac{2(\sum o_{j})}{2np-p}B^{\prime}+B^{\prime}\\ \mathbf{0}_{d+1}.\end{bmatrix}.$
@@
$\left\|\bm{\beta}^{*}-\frac{2}{2np-p}\sum_{j:o_{j}=1}\mathbf{x}_{n}^{\top}\bm{\beta}^{(K+1)}_{j}\right\|\leq$ $\left\|\bm{\beta}^{*}-\frac{1}{np-\frac{p}{2}}\sum_{j:o_{j}\in\mathcal{O}}\bm{\beta}^{(K+1)}_{j}\right\|+\left\|\frac{1}{np-\frac{p}{2}}\sum_{j:o_{j}\in\{o_{j}^{\prime}\}}\mathbf{x}_{n}^{\top}\bm{\beta}^{(K+1)}_{j}\right\|$
@@
Denote $\delta^{\prime}$ as $\delta^{\prime}=\mathbb{P}\{\neg\mathcal{E}_{1}\}+\mathbb{P}\{\neg\mathcal{E}_{2}(K\sqrt{np}(1-2\gamma/2)^{K})$, then the theorem follows.
Next, we establish the result concerning the label prediction loss presented in Theorem 5.2 through the following corollary:

### v2

@@ Appendix A Additional Related Works
Training Dynamics of Transformers.
There exist some works aiming to theoretically understand the ICL mechanism in Transformer through their training dynamics. Ahn et al., (2024); Mahankali et al., (2023); Zhang et al., (2023); Huang et al., (2023) investigate the dynamics of Transformers with a single attention layer and a single head for in-context linear regression tasks. Cui et al., (2024) prove that Transformers with multi-head attention layers outperform those with single-head attention. Cheng et al., (2023) show that local optimal solutions in Transformers can perform gradient descent in-context for non-linear functions. Kim and Suzuki, (2024) study the non-convex mean-field dynamics of Transformers, and Nichani et al., (2024) characterize the convergence rate for the training loss in learning a causal graph. Additionally, Chen et al., 2024a investigate the gradient flow in training multi-head single-layer Transformers for multi-task linear regression. Chen and Li, (2024) propose a supervised training algorithm for multi-head Transformers.
The training dynamics of Transformers for binary classification (Tarzanagh et al., 2023b, ; Tarzanagh et al., 2023a, ; Vasudeva et al.,, 2024; Li et al.,, 2023; Deora et al.,, 2023; Li et al., 2024a, ) and next-token prediction (NTP) (Tian et al., 2023a, ; Tian et al., 2023b, ; Li et al., 2024b, ; Huang et al.,, 2024) have also been studied recently.
@@ C.1 Transformer Structure
where the submatrix $\mathbf{W}_{1,\text{sub}}$ is defined as $\mathbf{W}_{1,\text{sub}}=\mathrm{diag}(\mathbf{0}_{(d+1)\times(d+1)},\mathbf{I}_{d\times d},0)$. Therefore, the output of the MLP layer is
@@
and $|\bm{\beta}_{2n+1}^{(k-1)}|\leq C_{\bm{\beta}}^{k}$. Then, we obtain
$|\bm{\beta}_{i}^{(k)}|$
$\leq|\bm{\beta}_{i}^{(k-1)}|+|\gamma^{(k-1)}|\|\mathbf{M}^{V}\|\left\|\frac{1}{i}[\mathbf{X}]_{1:\lfloor\frac{i+1}{2}\rfloor,:}^{\top}[\mathbf{X}]_{1:\lfloor\frac{i+1}{2}\rfloor,:}\right\||\bm{\beta}_{i}^{(k-1)}|+b_{\bm{\beta}}|\gamma^{(k-1)}|\|\mathbf{M}^{V}\|\left\|\frac{1}{i}[\mathbf{X}]_{1:\lfloor\frac{i+1}{2}\rfloor,:}\right\|$
@@
Equation C.9 arises from the fact that $\lfloor\frac{i+1}{2}\rfloor/i\leq 1$. Denoting $\gamma_{\max}=\max_{k}\|\gamma^{(k)}\|$ and letting $C_{\bm{\beta}}=1+b_{\mathbf{x}}^{2}d\gamma_{\max}\|\mathbf{M}^{V}\|+b_{\bm{\beta}}b_{\mathbf{x}}\gamma_{\max}\|\mathbf{M}^{V}\|$, we have
@@
$\sum_{u\in\{-1,+1\}}\sum_{j=1}^{i}\sigma\left(\Big{\langle}\mathbf{Q}^{(k)}_{u}\mathbf{h}_{i}^{k},\mathbf{K}^{(k)}_{u}\mathbf{h}_{j}^{k}\Big{\rangle}\right)\cdot\mathbf{V}_{u}^{(k)}\mathbf{h}^{(k)}_{j}$
$=\sum_{u\in\{-1,+1\}}\sum_{j=2}^{i-1}\sigma\left(\Big{\langle}\mathbf{Q}^{(k)}_{u}\mathbf{h}_{i}^{k},\mathbf{K}^{(k)}_{u}\mathbf{h}_{j}^{k}\Big{\rangle}\right)\cdot\mathbf{V}_{u}^{(k)}\mathbf{h}^{(k)}_{j}$
@@
$\sum_{u\in\{-2,+2\}}\sum_{j=1}^{i}\sigma\left(\Big{\langle}\mathbf{Q}^{(k)}_{u}\mathbf{h}_{i}^{k},\mathbf{K}^{(k)}_{u}\mathbf{h}_{j}^{k}\Big{\rangle}\right)\cdot\mathbf{V}_{u}^{(k)}\mathbf{h}^{(k)}_{j}$
$=\sum_{u\in\{-2,+2\}}\sum_{j=2}^{i-1}\sigma\left(\Big{\langle}\mathbf{Q}^{(k)}_{u}\mathbf{h}_{i}^{k},\mathbf{K}^{(k)}_{u}\mathbf{h}_{j}^{k}\Big{\rangle}\right)\cdot\mathbf{V}_{u}^{(k)}\mathbf{h}^{(k)}_{j}$
@@
$\mathbf{h}_{i}^{(k+1)}$ $=\mathrm{MLP}\left(\mathbf{h}_{i}^{(k)}+\frac{1}{i}\sum_{u\in\{\pm 1,\pm 2\}}\sum_{j=1}^{i}\sigma\left(\Big{\langle}\mathbf{Q}^{(k)}_{u}\mathbf{h}_{i}^{k},\mathbf{K}^{(k)}_{u}\mathbf{h}_{j}^{k}\Big{\rangle}\right)\cdot\mathbf{V}_{u}^{(k)}\mathbf{h}^{(k)}_{j}\right)$
$=\begin{bmatrix}[\mathbf{X}^{\top}]_{:,\lfloor\frac{i+1}{2}\rfloor}\\ 0\\ \bm{\beta}^{k+1}_{i}\\ 1\end{bmatrix},$
@@
$\mathbf{h}_{i}^{(k+1)}$ $=\mathrm{MLP}\left(\mathbf{h}_{i}^{(k+1)}+\frac{1}{i}\sum_{u\in\{\pm 1,\pm 2\}}\sum_{j=1}^{i}\sigma\left(\Big{\langle}\mathbf{Q}^{(k)}_{u}\mathbf{h}_{i}^{k},\mathbf{K}^{(k)}_{u}\mathbf{h}_{j}^{k}\Big{\rangle}\right)\cdot\mathbf{V}_{u}^{(k)}\mathbf{h}^{(k)}_{j}\right)$
$=\begin{bmatrix}[\mathbf{X}^{\top}]_{:,\lfloor\frac{i+1}{2}\rfloor}\\ 0\\ \bm{\beta}^{k+1}_{i}\\ 0\end{bmatrix}.$ (C.13)
@@
Combining with $c_{2}=\log\left(\frac{1}{\gamma^{(k)}\sigma_{\max}(2S-1)+\left|1-\gamma^{(k)}\sigma_{\min}\right|}\right)$, we obtain
@@
$\max_{i\neq j}\left|(\mathbf{D}_{:,i})^{\top}\mathbf{X}_{:,j}\right|\leq\sqrt{\frac{\log d-\log\delta}{Nc}}.$
@@
Next, we introduce a proposition as following.
@@
$\gamma\leq\min\left\{\frac{1}{\sigma_{\min}},\frac{1}{\max\{([\mathbf{D}_{n}]_{:,i})^{\top}\mathbf{X}_{1:n,i}\}}\right\}.$ (D.11)
Then, Inequality D.6 will hold and $1-\gamma\sigma_{\min}\geq 0$. Besides, we have
@@
we have $\gamma$ satisfies Equation D.11. Note that we assume $n\geq N_{0}$. Thus, we have
@@
$\|\bm{\beta}^{(k)}_{2n+1}-\bm{\beta}^{*}\|$
$\leq\|\bm{\beta}^{*}\|_{1}\Big{(}\gamma(2S-1)\sigma_{\max}+\left|1-\gamma\sigma_{\min}\right|\Big{)}^{k-1}$
@@ D.2 Proof of Remark 4
The output of the modified MLP layer is
@@
and the corresponding output of the self-attention layer satisfies
@@
Thus, the proof is complete. ∎
@@
$\frac{1}{i}\sum_{j=1}^{i}\left\langle\mathbf{Q}_{1}^{\mathrm{last}}\mathbf{h}_{i}^{(K+1)},\mathbf{K}_{1}^{\mathrm{last}}\mathbf{h}_{j}^{(K+1)}\right\rangle\cdot\mathbf{V}_{1}^{\mathrm{last}}\mathbf{h}_{j}^{(K+1)}$
$=\frac{1}{i}\sum_{j=1}^{i}\sigma\left(\mathbf{x}_{n}^{\top}\bm{\beta}^{(K+1)}_{j}-B^{\prime}\right)\cdot\begin{bmatrix}\mathbf{0}_{d}\\ \mathbbm{1}_{\{i\%2=1\}}\\ \mathbf{0}_{d+1}\end{bmatrix}$
@@
$\mathrm{MLP}\left(\frac{1}{i}\sum_{j=1}^{i}\left\langle\mathbf{Q}_{1}^{\mathrm{last}}\mathbf{h}_{i}^{(K+1)},\mathbf{K}_{1}^{\mathrm{last}}\mathbf{h}_{j}^{(K+1)}\right\rangle\cdot\mathbf{V}_{1}^{\mathrm{last}}\mathbf{h}_{j}^{(K+1)}\right)$
$=\begin{bmatrix}\mathbf{0}_{d}\\ \frac{2}{2np-p}\sum_{j:o_{j}=1}\mathbf{x}_{n}^{\top}\bm{\beta}^{(K+1)}_{j}-\frac{2(\sum o_{j})}{2np-p}B^{\prime}+B^{\prime}\\ \mathbf{0}_{d+1}.\end{bmatrix}.$
@@
$\left\|\bm{\beta}^{*}-\frac{2}{2np-p}\sum_{j:o_{j}=1}\mathbf{x}_{n}^{\top}\bm{\beta}^{(K+1)}_{j}\right\|$
$\leq\left\|\bm{\beta}^{*}-\frac{1}{np-\frac{p}{2}}\sum_{j:o_{j}\in\mathcal{O}}\bm{\beta}^{(K+1)}_{j}\right\|+\left\|\frac{1}{np-\frac{p}{2}}\sum_{j:o_{j}\in\{o_{j}^{\prime}\}}\mathbf{x}_{n}^{\top}\bm{\beta}^{(K+1)}_{j}\right\|$
@@
Then, the theorem follows by denoting $\delta^{\prime}$ as $\delta^{\prime}=\mathbb{P}\{\neg\mathcal{E}_{1}\}+\mathbb{P}\{\neg\mathcal{E}_{2}(K\sqrt{np}(1-2\gamma/2)^{K})$.
Finally, we establish the result concerning the label prediction loss presented in Theorem 5.2 through the following corollary: